Evaluate kronecker symbol (11/24)


1
Expert's answer
2021-02-24T06:32:46-0500

If "n=p_1...p_k," then "(\dfrac{d}{n})=(\dfrac{d}{p_1})...(\dfrac{d}{p_2})"


"24=2\cdot2\cdot2\cdot3"

"11\equiv3(mod\ 8)" Then "(\dfrac{11}{2})=-1"

"d(\dfrac{11}{3})" is the Legendre symbol: "d(\dfrac{11}{3})=-1"


"d(\dfrac{11}{24})=d(\dfrac{11}{2})d(\dfrac{11}{2})d(\dfrac{11}{2})d(\dfrac{11}{3})"

"=-1(-1)(-1)(-1)=1"

"d(\dfrac{11}{24})=1"